# thanks: current affiliation: IBM Quantum

### v1

@@ Acknowledgements.
Appendix B Creating robust logical sequences
As discussed above, we use the logical operators $\overline{XX}=XIXI$ and $\overline{YY}=IYIY\equiv YIYI$ (along with identity operators) to generate the LDD groups ${\mathcal{G}}_{{\text{LXY4}}}=\{I,\overline{XX},\overline{YY},\overline{ZZ}\}$, and ${\mathcal{G}}_{{\text{LXX}}}=\{I,\overline{XX}\}$. Cycling over the group elements creates the four-pulse sequences LXY4 and LXX shown in the red boxes of Fig. 1(d,e).
Since the main source of logical errors is $ZZ$ crosstalk, we optimize LDD so that it cancels such errors. Thus, in the case of LXX, we insert a stabilizer every other pulse-step to consistently apply a staggered sequence to all qubits. These sequences correspond to applying logical $X$ and $Y$ operators to both logical qubits.
To create the robust sequences, we mirror the original sequences (hence the eight-pulse sequences) and instead of $X$ and $Y$, we apply $\tilde{X}=R_{X}(-\pi)$ $\tilde{Y}=R_{Y}(-\pi)$, respectively, such that each sequence undergoes a robust pulse sequence at the physical level. We also apply staggered physical XY4 to each qubit as shown in Fig. 1(f). This sequence is also inherently robust at the physical level since each qubit receives an XY4 sequence, which is robust to pulse errors [62]. However, it is not a universal decoupling sequence at the logical level, as seen in the figure (i.e., it lacks the ability to decouple arbitrary single logical-qubit errors).
Appendix C The encoding/unencoding Lemma
Encoding of stabilizer codes is a familiar problem [91]. Here we present an explicit version of encoding circuits that appears to be simpler than other versions found in the literature [92, 93].
C.1 Lemma and proof
Lemma 1.
For any $[[n,k,d]]$ stabilizer code it is always possible to construct a unencoding circuit $V^{\dagger}$ such that $V^{\dagger}$ converts any group of canonical logical operators and a corresponding stabilizer group into single-qubit Pauli $X$ and $Z$ operators acting just on the first $k$ physical data qubits and into single-qubit Pauli $Z$ stabilizer operators on the remaining $n-k$ ancilla qubits.
Proof.
Our proof is constructive and relies on the basic Clifford group identities $\text{CNOT}_{i,j}(X_{i}X_{j})\text{CNOT}_{i,j}=X_{i}$ and $\text{CNOT}_{i,j}(Z_{i}Z_{j})\text{CNOT}_{i,j}=Z_{j}$, where $\text{CNOT}_{i,j}$ denotes the controlled-NOT gate with qubit $i$ as the control and qubit $j$ as the target. Note that this identity allows us to reduce the number of $X$ operators.
Consider first the case $k>0$.
  1. 1.
    Order the canonical logical operators and stabilizers as $\{\overline{X}_{1},\overline{Z}_{1},\dots,\overline{X}_{k},\overline{Z}_{k},S_{1},...,S_{n-k}\}$.
  2. 2.
    Ensure that $\overline{X}_{1}$ consists only of $X$ and $I$ physical operators: apply $H=(X+Z)/\sqrt{2}$ and $(X+Y)/\sqrt{2}$ to the qubits where $Z$ and $Y$, respectively, are present in the expression for $\overline{X}_{1}$.
  3. 3.
    Apply SWAP1,j if necessary to ensure $X_{1}$ is present in the expression for $\overline{X}_{1}$.
  4. 4.
    Apply $Z_{1}$ if necessary to cancel the sign if the sign was $-1$.
  5. 5.
    Apply CNOT1,j to eliminate all $X$ components of $\overline{X}_{1}$ except $X_{1}$.
  6. 6.
    Apply analogs of steps 2, 4, and 5 for $\overline{Z}_{1}$.
  7. 7.
    Now note that all the other operators contain $I_{1}$ (indeed, we have $\overline{X}_{1}=X_{1}$, $\overline{Z}_{1}=Z_{1}$, and all other operators commute with these two, hence they cannot contain $X_{1}$, $Z_{1}$, or $Y_{1}$), so the task is reduced to the case with $k:=k-1$ and $n:=n-1$; proceed recursively.
Next, consider the case $k=0$. We have stabilizer operators $S_{1},...,S_{n}$ on $n$ physical qubits and wish to transform them to $Z_{1},...,Z_{n}$.
  1. 1.
    Apply analogs of previous steps 2-5 to $S_{1}$ to transform it into $Z_{1}$.
  2. 2.
    Recursively call this procedure to transform the remaining $n-1$ components of the remaining $n-1$ stabilizers to $Z_{2}$, …, $Z_{n}$.
  3. 3.
    Apply CNOT1,j gates to cancel the $Z_{1}$ components of the remaining $n-1$ stabilizers (if they had one). Note that they could not have had $X_{1}$ or $Y_{1}$ components because otherwise they would not commute with $S_{1}$.
  4. 4.
    Apply $X_{j}$ to cancel the sign of the remaining stabilizers if necessary.
C.2 Illustration using the $[[4,2,2]]$ code
As an example, consider the $[[4,2,2]]$ code. The stabilizer is $\mathcal{S}=\langle X^{\otimes 4},Z^{\otimes 4}\rangle$, and we choose the canonical logical operators as $\mathcal{L}=\langle\overline{X}_{1}=X_{1}X_{4},\overline{X}_{2}=X_{1}X_{2},\overline{Z}_{1}=Z_{1}Z_{2},\overline{Z}_{2}=Z_{1}Z_{4}\rangle$. The corresponding code basis states are
|$\ket{\overline{00}}=\frac{1}{\sqrt{2}}\ket{0000}+\ket{1111}$ (9)
|$\ket{\overline{10}}=\frac{1}{\sqrt{2}}\ket{0110}+\ket{1001}$
|$\ket{\overline{01}}=\frac{1}{\sqrt{2}}\ket{0011}+\ket{1100}$
|$\ket{\overline{11}}=\frac{1}{\sqrt{2}}\ket{1010}+\ket{0101}.$
Following the steps in the proof of 1, we obtain the following series of transformations:
$\overline{X}_{1}$ $=X_{1}X_{4}\stackrel{{\scriptstyle\text{CNOT}_{14}}}{{\longmapsto}}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}$
$\overline{Z}_{1}$ $=Z_{1}Z_{2}\longmapsto Z_{1}Z_{2}\stackrel{{\scriptstyle\text{CNOT}_{21}}}{{\longmapsto}}Z_{1}{\longmapsto}Z_{1}{\longmapsto}Z_{1}{\longmapsto}Z_{1}$
$\overline{X}_{2}$ $=X_{1}X_{2}{\longmapsto}X_{1}X_{2}X_{4}{\longmapsto}X_{2}X_{4}\stackrel{{\scriptstyle\text{CNOT}_{24}}}{{\longmapsto}}X_{2}{\longmapsto}X_{2}{\longmapsto}X_{2}$
$\overline{Z}_{2}$ $=Z_{1}Z_{4}\longmapsto Z_{4}{\longmapsto}Z_{4}{\longmapsto}Z_{2}Z_{4}\stackrel{{\scriptstyle\text{CNOT}_{42}}}{{\longmapsto}}Z_{2}{\longmapsto}Z_{2}$
$S_{1}$ $=X^{\otimes 4}{\longmapsto}X_{1}X_{2}X_{3}{\longmapsto}X_{2}X_{3}{\longmapsto}X_{2}X_{3}X_{4}{\longmapsto}X_{3}X_{4}$
$\quad\stackrel{{\scriptstyle H_{3}H_{4}}}{{\longmapsto}}Z_{3}Z_{4}\stackrel{{\scriptstyle\text{CNOT}_{34}}}{{\longmapsto}}Z_{3}$
$S_{2}$ $=Z^{\otimes 4}{\longmapsto}Z_{2}Z_{3}Z_{4}{\longmapsto}Z_{2}Z_{3}Z_{4}{\longmapsto}Z_{3}Z_{4}{\longmapsto}Z_{3}Z_{4}$
$\quad{\longmapsto}Z_{3}Z_{4}{\longmapsto}Z_{4}{\longmapsto}Z_{4}\ ,$ (10)
where we have only indicated the transformations once, for the operator that induces the transformation per the proof of 1; e.g., $\text{CNOT}_{14}$ is induced by $\overline{X}_{1}$ but is applied to every operator in the first timestep, etc.
Figure 8: Performance of LDD on for an encoded |$\ket{\Phi_{+}}$ Bell state on ibm_marrakesh for nine set of qubits. (a) shows the logical errors and (b) shows the postselection results. This device features tunable couplers that significantly reduce the crosstalk. Nevertheless, accumulation of dephasing errors over time (up to $\approx 55\leavevmode\nobreak\ \mu$s) lead to the appearance of logical errors. The same LDD sequence we used in the main text successfully suppresses these errors, as shown in (a). Additionally, LDD with postselection outperforms postselection alone, as shown in (b).
Thus, the unencoding circuit, i.e., the circuit mapping every logical and stabilizer operator from the $[[4,2,2]]$ code to the corresponding operators of the trivial code, is
$V^{\dagger}=\text{CNOT}_{34}H_{4}H_{3}\text{CNOT}_{24}\text{CNOT}_{42}\text{CNOT}_{21}\text{CNOT}_{14},$ (11)
where $H_{i}$ is the Hadamard on the $i$th qubit.
After unencoding, we obtain the trivial code with $\mathcal{S}_{0}=\langle Z_{3},Z_{4}\rangle$ and $\mathcal{L}_{0}=\langle X_{1},X_{2},Z_{1},Z_{2}\rangle$, that divides the code into $k=2$ data qubits and $n-k=2$ ancillas.
In the trivial code, error correction is as follows: apply $IIXI$ when $IIZI=-1$ and $IIIX$ when $IIIZ=-1$. Applying $V^{\dagger}$ to these two recovery operators we find that to correct an undecoupled error in the original $[[4,2,2]]$ code, one should apply $IIIZ$ whenever $XXXX=-1$ and $XXXI$ (or $IIIX$, since they are related by a stabilizer element) when $ZZZZ=-1$.
@@ Appendix D Issues involving the implementation of logical state tomography using the $[[4,2,2]]$ code
Appendix E Logical dynamical decoupling with tunable couplers
Here we present additional experiments on ibm_marrakesh, which features tunable couplers [104] unlike the always-on $ZZ$ interaction in ibm_kyiv. Consequently, for this set of experiments, crosstalk is significantly reduced compared to the other set of results ($\lesssim 5$ kHz vs tens of kHz). Nevertheless, a combination of dephasing and residual crosstalk accumulates, leading to logical errors. Figure 8 illustrates the performance of LDD sequences on this QPU, averaged over 9 different qubit sets. Fig. 8(a) [equivalent to Fig. 3(b)] shows that without DD, logical errors remain unsuppressed. However, both logical DD sequences we employ successfully suppress these errors. Fig. 8(b) [equivalent to Fig. 5(a)] demonstrates that using only postselection, the fidelity averages to $\mathcal{F}\approx 56.23\%$, whereas combining postselection with LDD increases it to $\mathcal{F}\approx 91.73\%$.
Optimizing LDD for specific types of tunable couplers to maximize the interplay between DD and couplers is an avenue for future work.

### v2

@@ Appendix A Quantum codes and logical errors
Appendix C Construction of our $[[4,2,2]]$ code subgroup-NDD sequence
One of the limitations of our experimental setup is that physical $Z$ gates are unavailable: fixed-frequency transmons utilize virtual-$Z$ gates [96, 88]. Such gates are problematic for DD [88]. Here, now demonstrate that there are DD sequences satisfying the conditions of 1 which do not use any physical $Z$ pulses in the context of the $[[4,2,2]]$ code.
To construct such a sequence, we first choose generators of $\mathcal{L}$ that do not involve physical $Z$’s. A simple example is $\mathcal{L}=\langle i,XIIX,IIXX,IIYY,YIIY\rangle$. One could follow the construction below with this choice. However, in order to improve the ability of the DD sequence to suppress $ZZ$ crosstalk, we make a different choice. First, multiply the original generators of $\mathcal{L}$ (that is, $XIIX,IIXX,IIZZ,ZIIZ$) by the stabilizers $XXXX$, $XXXX$, $XXXX$, and $ZZZZ$, respectively, to obtain $\mathcal{L}^{\prime}=\langle i,IXXI,XXII,XXYY,IZZI\rangle$. Then, after some group operations, we can write the same group using a different set of generators containing no $Z$’s: $\mathcal{L}^{\prime}=\langle i,XXII,IIYY,IYIY,XIXI\rangle$. The corresponding NDD sequence consisting of two repetitions of $XIXI$, $IYIY$, $XIXI$, $IIYY$, $XIXI$, $IYIY$, $XIXI$, and $XXII$, which implements the decoupling group /{±$\mathcal{L}^{\prime}/\{\pm 1,\pm i\}$, and involves no $Z$ gates.
This sequence was obtained as follows: denote the generators of $\mathcal{L}$ (or $\mathcal{L}^{\prime}$) as $h_{1},h_{2},h_{3},h_{4}$ and choose any Gray code—a sequence of bitstrings $a_{0},\dots,a_{15}$ such that neighboring bitstrings differ in only a single digit. Then, pick $g_{j}=h^{a_{j}}$ (for $j=0,\dots,15$), where $a_{j}$ is interpreted as a multiindex (i.e., $g_{j}=\prod_{k=1}^{4}h_{k}^{(a_{j})_{k}}$). The pulses of the DD sequence are $g_{j}g_{j+1}$ for $j=0,\dots,15$.
@@ Appendix D Issues involving the implementation of logical state tomography using the $[[4,2,2]]$ code
Appendix E Creating robust logical sequences
As discussed in the main text, we use the logical operators $\overline{XX}=XIXI$ and $\overline{YY}=IYIY\equiv YIYI$ (along with identity operators) to generate the LDD groups ${\mathcal{G}}_{{\text{LXY4}}}=\{I,\overline{XX},\overline{YY},\overline{ZZ}\}$, and ${\mathcal{G}}_{{\text{LXX}}}=\{I,\overline{XX}\}$. Cycling over the group elements creates the basis for the four-pulse sequences shown in the red boxes of Fig. 1(d,e). However, we add two additional components to enhance crosstalk cancellation and robustness.
Since the main source of logical errors is $ZZ$ crosstalk, we modify LDD so that it cancels such errors. Thus, in the case of LXX, we insert a stabilizer every other pulse-step to consistently apply a staggered sequence to all qubits; in this manner, we move from LDD to NDD. These sequences correspond to applying logical $X$ and $Y$ operators to both logical qubits.
To create the robust sequences, we mirror the original sequences (hence the eight-pulse sequences) and instead of $X$ and $Y$, we apply $\tilde{X}=R_{x}(-\pi)$, $\tilde{Y}=R_{y}(-\pi)$, respectively, such that each sequence undergoes a robust pulse sequence at the physical level. We also apply staggered physical XY4 to each qubit as shown in Fig. 1(f). This sequence is also inherently robust at the physical level since each qubit receives an XY4 sequence, which is robust to pulse errors [62]. However, it is not a universal decoupling sequence at the logical level, as seen in the figure (i.e., it lacks the ability to decouple arbitrary single logical-qubit errors).
Figure 8: Performance of NDD on for an encoded |$\ket{\Phi_{+}}$ Bell state on ibm_marrakesh for a single set of qubits. (a) shows the logical errors and (b) shows the postselection results. This device features tunable couplers that significantly reduce the crosstalk. Nevertheless, accumulation of dephasing errors over time (up to $\approx 55\leavevmode\nobreak\ \mu$s) leads to the appearance of logical errors. The same NDD sequence we used in the main text successfully suppresses these errors, as shown in (a). Additionally, NDD with postselection outperforms postselection alone, as shown in (b).
Appendix F Normalizer dynamical decoupling with tunable couplers
Here we present additional experiments on ibm_marrakesh, which features tunable couplers [98] unlike the always-on $ZZ$ interaction in ibm_kyiv. Consequently, for this set of experiments, crosstalk is significantly reduced compared to the other set of results ($\lesssim 5$ kHz vs tens of kHz). Nevertheless, a combination of dephasing and residual crosstalk accumulates, leading to logical errors. Figure 8 illustrates the performance of NDD sequences on this QPU, averaged over $9$ different qubit sets. Figure 8(a) [equivalent to Fig. 3(b)] shows that without DD, logical errors remain unsuppressed. However, both NDD sequences we employ successfully suppress these errors. Figure 8(b) [equivalent to Fig. 5(a)] demonstrates that using only postselection, the fidelity averages to $\mathcal{F}\approx 56.23\%$, whereas combining postselection with NDD increases it to $\mathcal{F}\approx 91.73\%$.
Optimizing NDD for specific types of tunable couplers to maximize the interplay between DD and couplers is an avenue for future work.